The city of Suburbia received a proposal for plant watering from Water Can Gardening Company. The company would water all city plants for $469 each quarter, increasing the amount charged by $18 each subsequent quarter. Water Can Gardening Company requires a 10-year commitment.

The mayor has asked you to calculate the equivalent uniform quarterly cost for the plant water service to help with budgeting. Use a nominal annual interest rate of 6%, compounded quarterly, to make the calculation.
Answer #1

I = 6% = 6% / 4 = 1.5% per quarter

t = 10 yrs = 10 * 4 = 40 quarters

Equivalent quarterly cost = 469 + 18 * (A/G, 1.5%,40)

= 469 + 18 * 17.527729

= 784.50